# On the escape rate for intermittent maps with holes shrinking around the indifferent fixed point

### v1

@@ 3. The open systems and the escape rate
Let $m$ be the Lebesgue measure. Then, with $a_{N}\sim N^{-\frac{1}{s}}$, we obtain for $s\not=1$
$p_{N}\sim N^{-\frac{1}{s}-1}\quad\text{and}\quad\sum\limits_{k=1}^{N}\,k\,p_{k}\sim N^{-\frac{1}{s}+1}\sim m(H)^{1-s}\,$,
@@
$\gamma_{\mu}(H)\sim\left\{\begin{array}[]{ll}m(H)^{s}\,,&\text{if $s\not=1$;}\\[5.69046pt] \frac{m(H)}{-\log m(H)}\,,&\text{if $s=1$.}\end{array}\right$.
@@
(3.10) $\gamma_{\mu}(H)\sim\left\{\begin{array}[]{ll}const\cdot m(H)^{s}\,,&\text{if $s\not=1$;}\\[5.69046pt] const\cdot\frac{m(H)}{-\log m(H)}\,,&\text{if $s=1$,}\end{array}\right$.
@@
implies the result for $s=1$. ∎
@@
$\gamma_{\mu}(H_{\varepsilon})\asymp\left\{\begin{array}[]{ll}m(H_{\varepsilon})^{s}\,,&\text{if $s\not=1$;}\\[5.69046pt] \frac{m(H_{\varepsilon})}{-\log m(H_{\varepsilon})}\,,&\text{if $s=1$.}\end{array}\right$.
@@
Let’s consider first the case $s\not=1$. Then, there exists a constant $C>0$ such that for $\varepsilon$ small enough, hence $N$ big enough,
@@
This concludes the proof. ∎

### v2

@@ 3. The open systems and the escape rate
Let $m$ be the Lebesgue measure. Then, as $N\to\infty$, with $a_{N}\sim N^{-\frac{1}{s}}$, we obtain for $s\not=1$
$p_{N}\sim N^{-\frac{1}{s}-1}\quad\text{and}\quad\sum\limits_{k=1}^{N}\,k\,p_{k}\left\{\begin{array}[]{ll}=O(1)\,,&\text{for $s\in(0,1)$;}\\[5.69046pt] \sim N^{-\frac{1}{s}+1}\sim m(H)^{1-s}\,,&\text{for $s>1$;}\end{array}\right$.
@@
$\gamma_{\mu}(H)\sim\left\{\begin{array}[]{ll}m(H)\,,&\text{if $s\in(0,1)$;}\\[5.69046pt] \frac{m(H)}{-\log m(H)}\,,&\text{if $s=1$;}\\[5.69046pt] m(H)^{s}\,,&\text{if $s>1$.}\end{array}\right$.
@@
(3.10) $\gamma_{\mu}(H)\sim\left\{\begin{array}[]{ll}const\cdot m(H)\,,&\text{if $s\in(0,1)$;}\\[5.69046pt] const\cdot\frac{m(H)}{-\log m(H)}\,,&\text{if $s=1$;}\\[5.69046pt] const\cdot m(H)^{s}\,,&\text{if $s>1$.}\end{array}\right$.
@@
implies the result for $s=1$.
The case $s\in(0,1)$ follows from (3.11), since the denominator in $\gamma_{\mu}(H)$ is summable in this case. ∎
@@
$\gamma_{\mu}(H_{\varepsilon})\asymp\left\{\begin{array}[]{ll}m(H)\,,&\text{if $s\in(0,1)$;}\\[5.69046pt] \frac{m(H)}{-\log m(H)}\,,&\text{if $s=1$;}\\[5.69046pt] m(H)^{s}\,,&\text{if $s>1$.}\end{array}\right$.
@@
Let’s consider first the case $s>1$. Then, there exists a constant $C>0$ such that for $\varepsilon$ small enough, hence $N$ big enough,
@@
The same argument for $s\in(0,1)$ concludes the proof. ∎